OpenStudy (anonymous):

Challenge: How many 6-sided dice does it take to make rolling any 6 more likely than not?

OpenStudy (anonymous):

dont they all have the same chances?

OpenStudy (anonymous):

If you roll two dice, the chances for getting a six at all is greater than 1/6

OpenStudy (mathmate):

It's easier to consider the chances of NOT having even one six. For one die, P(1)=5/6 For two, P(2)=(5/6)^2 and for n dice, P(n)=(5/6)^n We want P(n)=(5/6)^n <=0.5 Take log and solve for n to get n=log(0.5)/log(5/6)=3.8 So with 4 dice, the probability of getting at least one 6 is more than half.

OpenStudy (anonymous):

Which is why Risk only gives you three. It all makes sense.

OpenStudy (mathmate):

:)